OpenStudy (kmullis6):

The table below represents a linear function f(x) and the equation represents a function g(x):

OpenStudy (kmullis6):

@ganeshie8

ganeshie8 (ganeshie8):

your main questiona nd the attached figure don't match could you double check

ganeshie8 (ganeshie8):

you're talking about turtles in the main question but i don't see any turtles in ur attached figure

OpenStudy (kmullis6):

Yea sorry, I accidentally posted the wrong picture I'd had saved.

OpenStudy (kmullis6):

But I changed it to the one it is now.

ganeshie8 (ganeshie8):

okay good

OpenStudy (kmullis6):

I know how to find slope, but I don't really get which one I'm supposed to use

ganeshie8 (ganeshie8):

Look at the table it is given the function in that table represents a line

ganeshie8 (ganeshie8):

can you find the slope of the function in that table ?

OpenStudy (kmullis6):

x: diff. is 1 f(x): diff. is 10

ganeshie8 (ganeshie8):

Yes!

ganeshie8 (ganeshie8):

so, slope of the function in table = ?

OpenStudy (kmullis6):

What do I say in place of what I put earlier?

ganeshie8 (ganeshie8):

slope = f(x) diff / x diff = 10 / 1 = 10

ganeshie8 (ganeshie8):

what about the slope of g(x) do you know how to find th slope of a line, given its equation ?

OpenStudy (kmullis6):

The slope of g(x) is 5. y = mx + b... m is the slope of the line

ganeshie8 (ganeshie8):

Excellent!

ganeshie8 (ganeshie8):

Now that you know the slopes of both the given functions, what is your answer for part A

OpenStudy (kmullis6):

The slope of x is 1, the slope of f(x) is 10, and the slope of g(x) is 5.

ganeshie8 (ganeshie8):

what do you mean by slope of x is 1 ? x is not a line you need a line before talking about slope

ganeshie8 (ganeshie8):

we saw that the slope of f(x) is 10, and the slope of g(x) is 5.

OpenStudy (kmullis6):

Oh, so what I said about it earlier didn't matter? Like it wasn't right?

ganeshie8 (ganeshie8):

` slope of f(x) is 10, and the slope of g(x) is 5. ` above part is correct

ganeshie8 (ganeshie8):

so which function has greater slope ?

OpenStudy (kmullis6):

f(x) does. Is this better? The slope of f(x) is 10, and the slope of g(x) is 5. In a table, you have to count the difference between the numbers, whereas in an equation, y = mx + b, m is the slope.

ganeshie8 (ganeshie8):

Looks good! always keep in mind, the more sentences you use to explain, the more your teacher likes you